There is 100 N of gravitational force between two objects. If these are moved so that they are now twice as far apart, what is the new gravitational force between them.
Group of answer choice

the is 100 N of gravitational force between two objects. If the masses of both objects are doubled, what is the new gravitational force between them?
Group of answer choices


Felix Baumgartner sky dives starting 23 miles above Earth’s surface. If he and his space suit together have a mass of 110 kg, what is acceleration (in m/s2) due to Earth’s gravity right where he first jumps? The mass of the Earth is 5.98 x 1024 kg and the radius of the Earth is 6,370 km. (1.61 km = 1 mi)
1
Expert's answer
2020-04-08T10:30:26-0400

a)


"F'=G\frac{mM}{(2r)^2}=G\frac{mM}{4(r)^2}=0.25F=0.25(100)=25\ N"

b)


"F'=G\frac{2m\cdot2M}{(r)^2}=4G\frac{mM}{(r)^2}=4F=4(100)=400\ N"

c)


"g=G\frac{M}{(R+h)^2}=6.67\cdot10^{-11}\frac{5.98 \cdot10^{24}}{(6.37
\cdot10^{6}+23(1610))^2}=\\=9.72\frac{m}{s^2}"